Wednesday, June 13, 2012

Day 16, 6/12/2012 -- Cooling

The cryostat continues to cool as it boils off helium at a copious rate. After 4 fills today (at ~150L of helium used apiece), the boiloff is now settling down -- the overnight fill should last for about 10-12 hours. The optics box was at about 65K in the evening; once it reaches about 40K the cooling rate increases and cools to near 10K relatively quickly. At that point, the fridges are cycled to cool the optics box and focal planes the rest of the way to 4K (and below).

This will be my last EBEX blog update for a while. I'm returning to Minnesota while Kate is coming here to help with detector work and calibration measurements for the next few weeks. I will take over Kate's job in MN of testing new detector wafers. I'll likely be back in Texas toward the end of June.

No pictures today, as everything in the high bay is pretty copacetic at the moment.

No comments:

Post a Comment

Note: Only a member of this blog may post a comment.